OpenStudy (anonymous):

group theory. Am I correct in saying (R,o) with x o y: lx . y| has no identity element?

OpenStudy (anonymous):

i was thinking because i can only think of 0 being the neutral identity but \( |x * 0| = |0 * x| = |x| \) if as it is over the reals, if x is -ve then -x does not equal |x|.

OpenStudy (loser66):

can you explain me why you do that?

OpenStudy (anonymous):

well the neural element is an element, say \(e\) such that \(a*e = e*a = a \)

OpenStudy (loser66):

by definition of neural element, you don't have that form for (R,0) because R*0 =0 it's not =R ---> it doesn't have identity.

OpenStudy (anonymous):

oops i meant 1 not 0, ie x*1

OpenStudy (loser66):

whatever, the problem is 0 , how can you get R when multiple R with 0?

OpenStudy (anonymous):

there is no multiplying by 0, thats what i mean it should be multiplying by 1.

OpenStudy (loser66):

I understand what you mean. the condition to be neural identity is the first term * the second term = the first term, ok?

OpenStudy (loser66):

if your couple is (3,1) then, it is "neural identity"

OpenStudy (anonymous):

no you have misunderstood

OpenStudy (loser66):

ok, explain me

OpenStudy (anonymous):

there should be one element, such that any other element 'multiplied' by that element equals itself

OpenStudy (loser66):

TedG, after explaining, you master your problem and you can answer it by yourself. hahahaa...

OpenStudy (anonymous):

now i say 'multiplied' because it in group theory it doesn't mean your normal multiplication. it is some defined binary operation.

OpenStudy (anonymous):

well i do sort of know i just wasnt sure haha

OpenStudy (loser66):

I guess!!! your group has more than 2 members than just R and 0

OpenStudy (loser66):

TedG, since your question is in group theory, I strongly recommend you make question at other site.

OpenStudy (anonymous):

well the set in question is (R, o) with x o y defined as |x . y|, where . is normal multiplication, and o is the defined binary operation.

OpenStudy (loser66):

want to have the link?

OpenStudy (anonymous):

no its fine i know the site.

OpenStudy (loser66):

ok, good luck